# For Better or For Worse? Learning Minimum Variance Features With Label Augmentation

### v1

@@ 3 Main Theoretical Results
Let $w^{*}$ be the unique minimizer of $\ell_{\lambda}(w)$ for $\lambda>0$. Then $\norm{w^{*}_{\mathcal{H}}}^{2}\geq\frac{1}{2}\norm{w^{*}}^{2}$. In other words, the optimal solution under weight decay has significant weight in the high variance dimensions $\mathcal{H}$.
Proof Sketch. We can orthogonally decompose the optimal solution $w^{*}$ in terms of the unit normal max-margin directions $u_{1}^{*}$ and $u_{2}^{*}$ separating $x_{\mathcal{L}}$ and $x_{\mathcal{H}}$ respectively. Then, using the fact that $y\left\langle u_{2}^{*},x_{\mathcal{H}}\right\rangle>y\left\langle u_{1}^{*},x_{\mathcal{L}}\right\rangle$ for all $(x,y)\in\mathrm{supp}(\pi)$ (from assumption three in Definition 2.1), we can claim that $w^{*}$ must have greater weight associated with $u_{2}^{*}$ than $u_{1}^{*}$, as otherwise we can decrease $\ell_{\lambda}(w^{*})$ by moving some weight from $u_{1}^{*}$ to $u_{2}^{*}$.
@@
On the other hand, as alluded to earlier, this phenomenon does not occur when considering the label smoothing and Mixup losses $\ell_{\mathrm{LS},\alpha}(w)$ and $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$. Indeed, we can show:
@@
Every minimizer $w^{*}$ of $\ell_{\mathrm{LS},\alpha}(w)$ for $\alpha>0$ satisfies $\norm{w^{*}_{\mathcal{H}}}=0$, so the optimal solutions under label smoothing have no weight in the high variance dimensions.
Proof Sketch. The optimal prediction at a label-smoothed point is either $21-\alpha/$ 2 or $\alpha/2$ (depending on whether the label is $1$ or $-1$ respectively). We can thus construct a predictor that is optimal at every point by choosing $w$ to have the appropriate correlation with $v_{1}^{*}$ and having no correlation with $x_{\mathcal{H}}$. That there are no equally good predictors that have correlation with $x_{\mathcal{H}}$ can be shown by appealing to convexity.
Analogously, we have the following result for Mixup:
@@
Every minimizer $w^{*}$ of $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ satisfies $\norm{w^{*}_{\mathcal{H}}}=0$ for any distribution $\mathcal{D}_{\lambda}$ with $\mathrm{supp}(\mathcal{D}_{\lambda})\subseteq[0,1]$ that is not a point mass on either 0 or 1.
Proof Sketch. The proof setup is similar to that of Theorem 3.3, i.e. we try to find a stationary point of $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$ that is zero in the high variance dimensions. The difficulty in the Mixup analysis is that it is not possible to directly construct a point-wise optimal predictor like in the proof of Theorem 3.3 due to the different possibilities for mixed points (and mixed labels), and consequently there is no obvious construction for a stationary point. However, we can analyze the limiting behavior of the Mixup gradient with respect to the weight correlation with the low variance dimensions and still show that such a stationary point must exist.
@@
Comparisons to existing results. We are not aware of any existing results on feature learning for label smoothing; prior theoretical work largely focuses on the relationship between label smoothing and learning under label noise, which is orthogonal to the perspective we take in this paper. Although feature learning results exist for Mixup (Chidambaram et al., 2023; Zou et al., 2023), the pre-existing results are constrained to the case of mixing using point mass distributions and only prove a separation between Mixup and unregularized ERM, whereas our results work for arbitrary mixing distributions (that don’t coincide with ERM) and also separate Mixup from ERM with weight decay. This greater generality comes at the cost of considering only linear models and linearly separable data, however, as both Chidambaram et al. (2023) and Zou et al. (2023) consider non-linearly separable data and the training dynamics of 2-layer neural networks.
@@
Additionally, our results can be viewed as generalizing the observations made by Chidambaram et al. (2023), which were that Mixup can learn multiple features in the data when doing so decreases the variance of the learned predictor. In our case, we directly show that Mixup will learn the low variance feature in the data as opposed to the high variance feature. Our results also do not contradict the observations of Zou et al. (2023), which were that Mixup can learn both a “common” feature and a “rare” feature in the data; in their setup the common/rare features are fixed (zero variance) per class and concatenated with high variance noise.
@@ 4 Experiments
(a) Weight Decay
(b) Label Smoothing
(c) Mixup
Figure 1: Final weight norm of logistic regression model in the high variance dimensions ($\norm{w_{\mathcal{H}}}$) for various hyperparameter settings on the synthetic data distribution of Section 4.1.
We now verify the predictions of our theoretical results in Section 3 across a variety of experimental setups. All experiments in this section were conducted on a single P100 GPU using PyTorch (Paszke et al., 2019) for model implementation. All reported results correspond to means over 5 training runs, and shaded regions in the figures correspond to 1 standard deviation bounds.
4.1 Synthetic Data
We first consider the case of training logistic regression models on a synthetic data distribution that exactly follows the assumptions of Definition 2.1.
Definition 4.1 (Synthetic Data).
We define a distribution $\pi_{\gamma}$ parameterized by $\gamma$ (where $0<\gamma\ll 1$) on $\mathbb{R}^{d}\times\{-1,1\}$ with $\pi_{\gamma,Y}(1)=1/2$ and $x\sim\pi_{\gamma,X\mid Y=y}$ satisfying:
  1. 1.
    First $\lfloor d/2\rfloor$ dimensions are constant but small. We have $x_{i}=\gamma y$ for $i\leq\lfloor d/2\rfloor$.
  2. 2.
    Last $d-\lfloor d/2\rfloor$ dimensions are high variance. We have $x_{i}\sim\mathrm{Uniform}([y,100y])$ for $i>\lfloor d/2\rfloor$.
In other words, we consider data where (conditional on the label) the first half of the dimensions are fixed (corresponding to $\mathcal{L}$) and the second half are i.i.d. high variance uniform (corresponding to $\mathcal{H}$). We sample $n=5000$ data points according to Definition 4.1 with $d=10$ and $\gamma=0.1$ and train logistic regression models across a range of settings for weight decay, label smoothing, and Mixup. [^1] The choice of $\gamma$ here and throughout the rest of this section is relatively arbitrary; we verified that our experimental phenomena are unchanged for other scales such as $10^{-2},10^{-3},\text{ and }10^{-4}$.
Namely, we consider 20 uniformly spaced values in $[0,0.1]$ for the weight decay $\lambda$ parameter and in $[0,0.75]$ for the label smoothing $\alpha$ parameter, where the upper bound for the label smoothing parameter space is obtained from the experiments of Müller et al. (2020). For Mixup, we fix the mixing distribution to be the canonical choice of $\mathrm{Beta}(\alpha,\alpha)$ introduced by Zhang et al. (2018) and consider 20 uniformly spaced $\alpha$ values in $[0,8]$ (with $\alpha=0$ corresponding to ERM), which effectively covers the range of Mixup hyperparameter values used in practice.
We train all models for 100 epochs using Adam with the standard hyperparameters of $\beta_{1}=0.9,\beta_{2}=0.999$, a learning rate of $5*10^{-3}$, and a batch size of 500. At the end of training, we compute $\norm{w_{\mathcal{H}}}$ (i.e. the norm of the weight vector in the last 5 dimensions) for each trained model. For each model setting, we report the mean and standard deviation of $\norm{w_{\mathcal{H}}}$ over 5 training runs in Figure 1.
As can be seen from the results, the weight decay models always have non-trivial values for $\norm{w_{\mathcal{H}}}$ (even for large values of $\lambda$), whereas the label smoothing and Mixup models very quickly converge to a $\norm{w_{\mathcal{H}}}$ of effectively zero as their respective hyperparameters move away from the ERM regime ($\alpha=0$ in both cases). This matches the behavior predicted by Theorems 3.1 to 3.4.
It is natural to ask to what extent these results depend on the fact that the low variance features in Definition 4.1 (the first half of the dimensions) actually have zero variance conditional on the label. As we mentioned in our theoretical results, it is sufficient to assume that the variance of the dimensions in $\mathcal{L}$ is simply much less than that of $\mathcal{H}$, and the fact that we consider zero variance is merely for simplifying purposes. We verify that this is true in practice by re-running the above experiments but modifying the dimensions in $\mathcal{L}$ to have non-zero variance by adding uniformly distributed noise on the scale of $\gamma/10$. The results remain unchanged; the corresponding plots can be found in Section B of the Appendix.
4.2 Binary Image Classification
@@
Although the experiments of Section 4.1 constitute a direct verification of our theory, in practice the low and high variance features are typically not as obviously partitioned as in Definition 4.1 and the metric we care about is usually generalization error as opposed to the strength of correlation with some part of the data. As such, we now relax our strict adherence to the assumptions of Definition 2.1 and consider binary classification on image data.
To do so, we reduce the standard benchmarks of CIFAR-10 and CIFAR-100 (Krizhevsky, 2009) to binary classification tasks by maintaining only the first two classes (“airplane” and “automobile” for CIFAR-10, “apple” and “aquarium fish” for CIFAR-100) and assigning them the labels of -1 and 1 respectively. This leads to 10000 training examples and 2000 test examples in the case of CIFAR-10, and 1000 training examples and 200 test examples in the case of CIFAR-100.
We preprocess the training data to have mean zero and variance one along every image channel (and apply the same transform to the test data), and then modify the training data such that the top left value (corresponding to the “R” channel of the top left pixel in the original image) in the tensor representation of the training input is replaced by $y\gamma$, where $\gamma=0.1$ as in Section 4.1. This ensures that the training data is separable in the first dimension of the data, although unlike in the case of Definition 4.1 we make no modifications to the other dimensions of the data to ensure separability outside the first dimension. We leave the test data unchanged - our goal is to determine whether models trained on the modified training data can learn more than just the single identifying dimension.
We follow the exact same logistic regression training setup as in Section 4.1, and sweep over the same hyperparameter values for weight decay, label smoothing, and Mixup. We report the mean test error along with one standard deviation error bounds over five training runs for each model setting in Figure 2.
The results mirror our observations in Section 4.1: both the label smoothing and Mixup models have high test error for all settings (in this case, ERM performs poorly as well) while weight decay achieves a significantly lower test error for all $\lambda>0$. The higher variance in the reported CIFAR-100 test errors can almost certainly be attributed to the fact that our binary version of CIFAR-100 consists of only 1000 training points and 200 points.
@@
Our results correspond once again to label smoothing and Mixup learning to use only the zero variance, identifying dimension of the input, as we expect from our theory. Indeed, we verify this in Figure 3 where we plot the ratio $\norm{w_{1}}/\norm{w_{[d]\setminus\{1\}}}$ (i.e. the ratio between the norm of the trained model weight vector in the first dimension and the remaining dimensions). We observe that for both CIFAR-10 and CIFAR-100, across all the non-ERM hyperparameter settings (refer to Section 4.1 for the range of hyperparameters) that both label smoothing and Mixup place much greater weight in the first dimension than weight decay.
As with the experiments of Section 4.1, we again check that the results here do not rely on the fact that the identifying dimension introduced into the training data has zero variance conditional on the label. As before, introducing uniformly distributed noise on the scale of $\gamma/10$ to the identifying dimension does not change our results, and the corresponding plots can again be found in Section B of the Appendix.
4.3 Multi-Class Image Classification
(a) Weight Decay ($\lambda=0.01$)
(b) Label Smoothing ($\alpha=0.1$)
(c) Mixup ($\alpha=1$)
Figure 4: ResNet-18 test error at the end of each epoch of training on the modified versions of CIFAR-10 and CIFAR-100 from Section 4.3, using fixed settings for weight decay, label smoothing, and Mixup.
Thus far we have considered training logistic regression models to stay within the confines of our theory, but we now investigate whether our observations extend to the more practical setting of training nonlinear models (i.e. deep neural networks) on multi-class image classification data.
As in Section 4.2, we consider CIFAR-10 and CIFAR-100, but this time maintain all of the original classes in both datasets. We once again preprocess the training data to have mean zero and variance one, and apply this transform to the test data as well. Unfortunately, unlike in Section 4.2, introducing a zero variance, identifying feature into each class in the training data would now require modifying a more significant portion of each image (for example, a $10\times 10$ patch for CIFAR-100 to identify each class).
Figure 5: Visualization of data modification applied to CIFAR-10/CIFAR-100 in Section 4.3.
So as to preserve most of the original image content, as well as to have a more direct comparison to prior work, we choose to adopt a similar experimental setup to that of Chidambaram et al. (2023). Namely, we assign a random vector sampled from $\mathcal{N}(0,\gamma^{2}\mathrm{Id}_{d})$ (where $d$ is the dimension of the data) to each class and concatenate this class vector along the channels dimension to every training data point for that class. A visualization of this transformation is shown in Figure 5. For the test data points, we concatenate an all-zero vector of the same size so as to preserve the dimensionality across train and test, as was done by Chidambaram et al. (2023). We take $\gamma=0.1/\sqrt{d}$, so as to account for the dimensionality of the concatenated feature.
For both datasets, we train ResNet-18 models (He et al., 2015) for 100 epochs using Adam with the standard hyperparameters of $\beta_{1}=0.9,\beta_{2}=0.999$, a learning rate of $10^{-3}$, and a batch size of 500. Due to the higher compute cost of training these models compared to the logistic regression models of Sections 4.1 and 4.2, we do not perform a full hyperparameter sweep but instead use known well-performing settings for weight decay, label smoothing, and Mixup. Namely, we use $\lambda=0.01$ for weight decay, $\alpha=0.1$ for label smoothing (Müller et al., 2020), and $\alpha=1$ in $\mathrm{Beta}(\alpha,\alpha)$ for Mixup (Zhang et al., 2018). Since we are now dealing with nonconvex training dynamics, we report the mean test error over five runs at the end of each epoch with one standard deviation error bounds for each model setting in Figure 4.
As can be seen from the plots, both label smoothing and Mixup initially decrease the test error in the initial stages of training, but then overfit to the concatenated zero variance features. On the other hand, the test error for weight decay decreases over the course of training until effectively converging. This once again does not contradict what was observed by Chidambaram et al. (2023), as in their case their concatenated features were weighted alongside with the original image data so that variance could be decreased by learning both.
We also considered the case of low variance concatenated features as opposed to zero variance concatenated features in the training data. Similar to the results of Sections 4.1 and 4.2, the label smoothing and Mixup overfitting phenomenons persist even when we introduce variance to the concatenated features. However, we found that unless the introduced variance is much smaller (i.e. multiple orders of magnitude) than the concatenated feature scale $\gamma$, weight decay also overfits to the concatenated features. This brittleness is possibly an artifact of the chosen experimental setup and/or hyperparameters, but due to compute constraints we did not exhaustively explore this behavior and instead leave it as an open direction for follow-up work extending our results to non-linear/non-convex settings.

### v2

@@ 3.1 Linear Binary Classification
Let $w^{*}$ be the unique minimizer of $\ell_{\beta}(w)$ for $\beta>0$ under $\pi$ satisfying Assumption 3.2. Then $\norm{w^{*}_{\mathcal{H}}}^{2}\geq\frac{1}{2}\norm{w^{*}}^{2}$.
Proof Sketch. We can orthogonally decompose the optimal solution $w^{*}$ in terms of unit normal directions $u^{*}\in\mathcal{L}$ and $v^{*}\in\mathcal{H}$. We show that in this decomposition it must be the case that $y\left\langle v^{*},x\right\rangle>y\left\langle u^{*},x\right\rangle$ for $\pi$-a.e. $(x,y)$, and then we can claim that $w^{*}$ must have greater weight associated with $v^{*}$ than $u^{*}$, as otherwise we can decrease $\ell_{\beta}(w^{*})$ by moving weight from $u^{*}$ to $v^{*}$.
@@
On the other hand, as alluded to earlier, this phenomenon does not occur when considering the label smoothing and Mixup losses $\ell_{\mathrm{LS},\alpha}$ and $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$. We will show that for both label smoothing and Mixup, the optimal solutions have arbitrarily small correlation with the dimensions in $\mathcal{H}$ as the distribution of $YX$ concentrates. For these results, we require the following assumptions (but no longer need Assumption 3.2).
Assumption 3.5.
We assume that $\mathbb{E}[YX_{\mathcal{L}}]\neq 0$ and $\Sigma_{YX,\mathcal{H}}\succ\rho\mathrm{Id}$ for some $\rho>0$. Here $\Sigma_{YX,\mathcal{H}}$ denotes the covariance matrix of $\Sigma_{YX}$ restricted to those rows and columns in $\mathcal{H}$.
The first part of Assumption 3.5 just ensures it is possible to obtain a good solution using the dimensions in $\mathcal{L}$ while the second part codifies the idea of $\mathcal{H}$ being a high (at least non-zero) variance feature. Observe that we have made no direct separability assumptions on the data, although it is true that as $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$ the class-conditional supports are guaranteed to be linearly separable in $\mathcal{L}$. We now state both results.
@@
For $\alpha\in(0,1)$ and $\pi$ satisfying Assumption 3.5, every minimizer $w^{*}$ of $\ell_{\mathrm{LS},\alpha}(w)$ satisfies $\norm{w^{*}_{\mathcal{H}}}<O(\norm{\Sigma_{YX,\mathcal{L}}})$.
Proof Sketch. We show that $\ell_{\mathrm{LS},\alpha}(w)$ is strongly convex in $w^{\top}YX$. We then use Jensen’s inequality to get a lower bound on the loss in terms of this quantity, and show that this lower bound is achievable using only $w\in\mathcal{L}$ as the variance of $YX_{\mathcal{L}}$ decreases. Then using a lower bound on the Jensen gap (Lemma A.1) and Assumption 3.5, we show that any solution that is sufficiently non-zero in the $\mathcal{H}$ dimensions remains bounded away from this lower bound.
@@
For any symmetric $\mathcal{D}_{\lambda}$ that is not a point mass on 00 or $1$ and $\pi$ satisfying Assumption 3.5, every minimizer $w^{*}$ of $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(w)$ satisfies $\norm{w^{*}_{\mathcal{H}}}<O(\norm{\Sigma_{YX,\mathcal{L}}})$.
Proof Sketch. The proof follows the same outline as the label smoothing result, but there are now additional difficulties due to mixing inputs. In particular, we need to condition on $\lambda$ and then show strong convexity of the conditional loss in terms of $w^{\top}Z_{\lambda}$. We can then get a lower bound on the overall loss as an expectation with respect to $\lambda$ of this conditional loss. However, it is no longer immediate that there exists a stationary point $w\in\mathcal{L}$ that minimizes this lower bound. We prove the existence of such a stationary point by considering limiting behavior of the gradient of the loss as $w_{\mathcal{L}}$ tends to $\infty$ or $\infty$- ∞ in each component. After establishing a stationary point, we can show as with the label smoothing result that as $\norm{\Sigma_{YX,\mathcal{L}}}\to 0$ we can get a solution with $w\in\mathcal{L}$ that achieves the lower bound while solutions with $w_{\mathcal{H}}$ non-zero are bounded away from the lower bound.
3.2 General Multi-Class Classification
Our results so far have demonstrated a separation between standard training (ERM + weight decay) and label augmentation (label smoothing and Mixup) in the linear binary classification setting, without explicitly having to assume linear separability. The benefit of the linear binary case is that in this case the learning problems are convex in the weight vector $w$, so we can directly discuss properties of the optimal solutions instead of worrying about optimization dynamics.
Of course, this is no longer true when we pass to nonlinear models, and standard model choices (e.g. neural networks) make it so that it is no longer easy to prove that a model has “learned” either $X_{\mathcal{L}}$ or $X_{\mathcal{H}}$ without explicitly analyzing some choice of optimization algorithm (i.e. gradient descent). However, our observations do translate to the outputs of any model.
It is obvious that the model output variance should go to zero for any model that achieves the global optimum of the label augmentation losses we have discussed; indeed, this just corresponds to predicting the correct labels ($\alpha,1-\alpha$ with label smoothing and $\lambda,1-\lambda$ for every $\lambda$ for Mixup) with probability 1. On the other hand, it is less clear that we can make a quantitative statement regarding how bad the loss could be given a certain amount of variance in the model output.
By lifting the techniques of the previous subsection, we can prove such quantitative results for both label smoothing and Mixup in the general multi-class setting. The general data distribution we consider for these results is as follows.
Definition 3.8.
[Multi-Class Data Distribution] We consider $\pi$ to be a distribution supported on $B\times[k]$ where $B$ is a compact subset of $\mathbb{R}^{d}$ and $k>2$. We assume only that $\pi$ satisfies $\pi_{Y}(y)>0$ for every $y$ and $\Sigma_{X}$ is positive definite.
We make virtually no assumptions on $\pi$ for these results because we will only be proving general lower bounds, which are weaker than the claims regarding the optimal solutions of the previous subsection. For $\pi$ as in Definition 3.8, the label smoothing and Mixup results are as follows.
Proposition 3.9.
For $\alpha>0$ and any $g:\mathbb{R}^{d}\to\Delta^{k-1}$, letting $\mathrm{OPT}_{\mathrm{LS},\alpha}$ denote the minimum of $\ell_{\mathrm{LS},\alpha}$, we have for a universal constant $C>0$:
$\ell_{\mathrm{LS},\alpha}(g)\geq\mathrm{OPT}_{\mathrm{LS},\alpha}+C\sum_{y=1}^{k}\pi_{Y}(y)\mathrm{Var}\left(g(X)\mid Y=y\right).$ (3.1)
Proposition 3.10.
For any $\mathcal{D}_{\lambda}$ that is not a point mass on 00 or $1$ and any $g:\mathbb{R}^{d}\to\Delta^{k-1}$, letting $\mathrm{OPT}_{\mathrm{MIX},\mathcal{D}_{\lambda}}$ denote the minimum of $\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}$, we have for a universal constant $C>0$ that:
$\ell_{\mathrm{MIX},\mathcal{D}_{\lambda}}(g)\geq\mathrm{OPT}_{\mathrm{MIX},\mathcal{D}_{\lambda}}+C\sum_{y_{1}=1}^{k}\sum_{y_{2}=1}^{k}\pi_{Y}(y_{1})\pi_{Y}(y_{2})\mathbb{E}_{\lambda}\left[\mathrm{Var}\left(g(Z_{\lambda})\mid y_{1},y_{2},\lambda\right)\right].$ (3.2)
The proofs for both Propositions 3.9 and 3.10 follow the same structure; we show that after appropriate conditioning, both losses can be broken up into a sum of strongly convex conditional losses. The lower bounds in both results show that in order to make progress with respect to the label smoothing or Mixup losses, a training algorithm needs to push model outputs to be low variance.
Remark 3.11.
We do not prove an analogous result to Theorem 3.3 in the general setting because Propositions 3.9 and 3.10 operate directly on the model outputs $g(X)$, and it is not clear how to translate a weight norm constraint to this setting without explicitly parameterizing $g$. Intuitively, however, with a weight norm constraint on $g$ it may no longer be optimal to have zero variance predictions, since such predictions may require very large weights depending on the data distribution.
@@ 3.2 General Multi-Class Classification
Comparisons to existing results. We are not aware of any existing results on feature learning for label smoothing; prior theoretical work largely focuses on the relationship between label smoothing and learning under label noise, which is orthogonal to the perspective we take in this paper. Although feature learning results exist for Mixup (Chidambaram et al., 2023; Zou et al., 2023), the pre-existing results are constrained to the case of mixing using point mass distributions and only prove a separation between Mixup and unregularized ERM, whereas our results work for arbitrary symmetric mixing distributions (that don’t coincide with ERM) and also separate Mixup from ERM with weight decay. This greater generality comes at the cost of considering only linear models for our feature learning results; both Chidambaram et al. (2023) and Zou et al. (2023) consider the training dynamics of 2-layer neural networks on non-separable data.
@@
Additionally, our results can be viewed as generalizing the observations made by Chidambaram et al. (2023), which were that Mixup can learn multiple features in the data when doing so decreases the variance of the learned predictor. In our case, we directly show that Mixup will have much larger correlation with the low variance feature in the data as opposed to the high variance feature. Our results also do not contradict the observations of Zou et al. (2023), which were that Mixup can learn both a “common” feature and a “rare” feature in the data; in their setup the common/rare features are fixed (zero variance) per class and concatenated with high variance noise.
@@ 4 Experiments
We now verify the predictions of our theoretical results in Section 3 across both binary and multi-class experimental setups. All experiments in this section were conducted on a single P100 GPU using PyTorch (Paszke et al., 2019) for model implementation. All reported results correspond to means over 5 training runs, and shaded regions in the figures correspond to 1 standard deviation bounds.
4.1 Spurious Pixels in Binary Classification
@@
We first consider the case of logistic regression trained on data that has trivial spurious correlations that do not appear in the test data. For our binary classification task, we consider reductions of the standard CIFAR-10 and CIFAR-100 (Krizhevsky, 2009) benchmarks to binary classification by fixing two classes from each dataset as the positive and negative classes, and replacing their original labels with the labels $+1$ and $-1$ respectively. This leads to 10000 training examples and 2000 test examples in the case of CIFAR-10, and 1000 training examples and 200 test examples in the case of CIFAR-100. Our experiments are not sensitive to the choice of binary reduction, and for the experiments in this section we fix the $-1$ class to be class 00 from the original data and the $+1$ class to be class $1$ from the original data (“airplane” and “automobile” for CIFAR-10, “apple” and “aquarium fish” for CIFAR-100).
We preprocess the training data to have mean zero and variance one along every image channel (and apply the same transform to the test data). To introduce spurious correlations into the training data, we modify the data such that the first value in the tensor representation of each training input (corresponding to the “R” channel of the top left pixel in the original image) is replaced by $y\gamma$ with $\gamma=0.1$ (the choice of $\gamma$ here is not important as long as it is small, we verified the results for $\gamma=10^{-5}$ up to $\gamma=0.1$). This ensures that the training data is trivially linearly separable in the first dimension of the data, but learning this first dimension requires having larger weight norm due to the scaling by $\gamma$. We leave the test data unchanged - our goal is to determine whether models trained on the modified training data can learn more than just the single identifying dimension.
We then train logistic regression models to convergence (100 epochs) on both the reduced CIFAR-10 and CIFAR-100 tasks across a range of settings for weight decay, label smoothing, and Mixup. We consider 20 uniformly spaced values in $[0,0.1]$ for the weight decay $\lambda$ parameter and in $[0,0.75]$ for the label smoothing $\alpha$ parameter, where the upper bound for the label smoothing parameter is obtained from the experiments of Müller et al. (2020). For Mixup, we fix the mixing distribution to be the canonical choice of $\mathrm{Beta}(\alpha,\alpha)$ introduced by Zhang et al. (2018) and consider 20 uniformly spaced $\alpha$ values in $[0,8]$ (with $\alpha=0$ corresponding to ERM), which effectively covers the range of Mixup hyperparameter values used in practice. We train all models using Adam with the hyperparameters of $\beta_{1}=0.9,\beta_{2}=0.999$, a learning rate of $5*10^{-3}$, and a batch size of 500.
After training, we report the mean test error along with one standard deviation error bounds over five training runs for each model setting in Figure 1. As can be seen from the results, both the label smoothing and Mixup models have high test error for all settings while weight decay achieves a significantly lower test error for all $\lambda>0$. The higher variance in the reported CIFAR-100 test errors can almost certainly be attributed to the fact that our binary version of CIFAR-100 consists of only 1000 training points and 200 points.
@@
Our results correspond to label smoothing and Mixup learning to use only the spurious, identifying dimension of the training inputs, as we expect from our theory since this dimension has zero conditional variance. Indeed, we verify this fact empirically in Figure 2, where we plot the ratio $\norm{w_{1}}/\norm{w_{[d]\setminus\{1\}}}$ (i.e. the ratio between the norm of the trained model weight vector in the first dimension and the remaining dimensions) for each of the trained models.
4.2 Background Shift in Multi-Class Classification
(a) Shaded MNIST
(b) Model Performance
Figure 3: Final model test errors over our hyperparameter sweep for the shaded MNIST dataset of Section 4.2, alongside a visualization of samples from the dataset.
Although the experiments of Section 4.1 cleanly verify our theory, the spurious correlation introduced in the training data is effectively adversarial in nature. We now consider the case of a spurious correlation of greater practical interest: backgrounds that are correlated with the correct class label.
In particular, we consider a very similar setup to the one used by Arjovsky et al. (2020) to motivate the influential invariant risk minimization framework: namely, we construct a shaded version of the MNIST dataset in which the background pixels for each class are replaced with different low intensity colors corresponding to the class labels. Unlike Arjovsky et al. (2020), however, we maintain all 10 classes since our theory from Section 3.2 suggests that our observations should generalize to the multi-class setting. The colors used for each class background are permuted between the train and the test data, so that a model that learns to predict only using background pixels cannot achieve good test accuracy.
The only constraints we place on the background colors are that they are distinct across classes and that their intensities (i.e. values in RGB space) are small (it suffices to consider values bounded by 16); this makes it difficult to visually distinguish the background colors (hence "shaded MNIST") even though they are obviously different in RGB space. The color intensity constraint is not necessary for the failure of label smoothing and Mixup, but is necessary for weight decay to succeed since it ensures that the higher variance feature (the actual digit in the foreground) is generally more separable (due to larger pixel intensities). Note that although there is no variance in the background color conditional on a class, there is significant variance in the actual pixels per class as the locations of the background pixels change across data points.
For our model setup, we also follow Arjovsky et al. (2020) and consider a simple 2-layer feedforward neural network with ReLU activations and a hidden layer size of $2048$, as this is sufficient to achieve good performance on MNIST and is a tiny enough model that we can efficiently do the same hyperparameter sweep from Section 4.1. We follow the same training procedure as Section 4.1, training until the training loss roughly converges (only 20 epochs are necessary).
Test error results across the different hyperparameter settings for weight decay, label smoothing, and Mixup are shown in Figure 3. We observe the same phenomena as before, even in this non-trivial spurious correlation setting: both the label smoothing and Mixup models have high test error for all settings while weight decay achieves a significantly lower test error for all $\lambda>0$.